# Indoor Air Quality Dataset with Activities of Daily Living in Low to Middle-income Communities

### v1

@@ Abstract
In recent years, indoor air pollution has posed a significant threat to our society, claiming over 3.2 million lives annually. Developing nations, such as India, are most affected since lack of knowledge, inadequate regulation, and outdoor air pollution lead to severe daily exposure to pollutants. However, only a limited number of studies have attempted to understand how indoor air pollution affects developing countries like India. To address this gap, we present spatiotemporal measurements of air quality from 30 indoor sites over six months during summer and winter seasons. The sites are geographically located across four regions of type: rural, suburban, and urban, covering the typical low to middle-income population in India. The dataset [^1] https://github.com/prasenjit52282/dalton-dataset (Access:July 19, 2024) contains various types of indoor environments (e.g., studio apartments, classrooms, research laboratories, food canteens, and residential households), and can provide the basis for data-driven learning model research aimed at coping with unique pollution patterns in developing countries. This unique dataset demands advanced data cleaning and imputation techniques for handling missing data due to power failure or network outages during data collection. Furthermore, through a simple speech-to-text application, we provide real-time indoor activity labels annotated by occupants. Therefore, environmentalists and ML enthusiasts can utilize this dataset to understand the complex patterns of the pollutants under different indoor activities, identify recurring sources of pollution, forecast exposure, improve floor plans and room structures of modern indoor designs, develop pollution-aware recommender systems, etc.
@@ 3.2 Annotating Activities of Daily Living
3.3 Measurement Sites & Demographics of the Participants
@@
In total, 46 occupants actively participated in the study by providing the activity and event context of their indoor space using the vocalAnnot application as discussed in Sec.  3.2. The participants include college students, university staff, professors, homemakers, canteen owners, etc. Among them, 27 were male, and 19 were female. Table 2 depicts the female participation ratio and the socioeconomic background of the occupants. The majority of the participants are between 20 and 40 years old. Thus, they are familiar with using Android apps during their everyday routines. Nevertheless, the youngest is a three-year-old child, while the oldest is an 84-year-old adult. Individuals, either too young or too old, do not actively participate in this study, and other family members report their indoor activities.
@@ 3.4 License and Consent
Figure 4: Split air conditioners (i.e., marked with S) suffer from compromised ventilation compared to the legacy window air conditioners (i.e., marked with W) simply to improve power efficiency. The figure clearly depicts the accumulation of VOC and CO2 over time when split AC is on. Meanwhile, we observe consistent ventilation when the window AC is on, achieving healthy air quality with time. The relative humidity is slightly higher for window AC (i.e., within the comfort range, 30–50%).
@@ 5 Limitations and Conclusion
To conclude, in this work, we have collected a cross-sectional indoor air quality dataset from low to middle-income households and indoor sites in India. Unlike publicly available data, our dataset captures the indoor pollution dynamics of developing countries. Data was collected over six months across 30 indoor sites, including research labs, canteens, classrooms, studio apartments, and residential properties, with 46 occupants providing annotations on indoor activities. Additionally, we include the geometry of the measurement sites, which can help understand the impact of floor plans on pollutant spread. We believe that this dataset can aid environmentalists, architects, and ML researchers in improving indoor designs, addressing pollution challenges, and developing models for predicting indoor pollution patterns.
@@ A.2 Composition
The dataset presents spatiotemporal air quality measurements from 30 indoor sites (e.g., studio apartments, classrooms, research laboratories, food canteens, and residential households) over six months during the summer and winter seasons (89.1M samples, totaling 13646 hours of air quality data and 3960 activity annotations from 46 participants).
@@ A.5 Preprocessing/cleaning/labeling
The dataset is cleaned and organized with the processing pipeline shown in Figure 8. First, the replicas from the log servers are merged, and any duplicates are dropped. Second, the sensor readings are validated based on the reported range of each sensor while marking any missing data segments. Next, missing values are imputed with the nearest fill method if the segment is less than 15 mins, zero otherwise. Further, the data is smoothed with a 60-second rolling window to reduce noise. Finally, the change-points in the sensor readings are marked to denote potential pollution events in the indoor space. After preprocessing, three new columns are computed from the sensor readings, as shown in the figure. The utility of the derived columns is as follows:
@@ A.8 Maintenance
Figure 9: Daily indoor trends by week and month. We observe higher concentrations of VOC and CO2 during the winter months.
@@ Appendix B More Dataset Information
Site ID Devices #Devices Site Area (Sqft) Floor Plan #F/#M Duration (Hrs) #Samples Annot Participants
@@ 95 Kitchen
A1 101 Study Desk 1 150 1/0 146 226888 P28
A2 105 Study Desk 1 150 0/1 289 193557 P29
A3 109 Study Desk 1 180 0/1 344 1098827 P30
A4 120 Study Desk 1 150 1/0 125 384975 P31
A5 121 Study Desk 1 150 1/0 1 77 P32
A6 122 Study Desk 1 100 0/1 51 154398 P33
A7 113 On bed 1 150 0/1 55 54741 P34
A8 151 Study Desk 1 150 0/1 60 189141 P35
@@ 52 Desk2
R2 102 Desk Lab 319 1 320 2/2 367 1161570 P43
R3 108 Left Desk 1 616 0/1 243 750745 P44
@@ 71 Desk1
F1 23 Kitchen 1 150 2/0 450 631193 P46
F2 24 Kitchen 1 150 450 631193
C1 25 Teacher Desk 1 500 333 590272
C2 26 Teacher Desk 1 500 53 158256

### v2

@@ Abstract
In recent years, indoor air pollution has posed a significant threat to our society, claiming over 3.2 million lives annually. Developing nations, such as India, are most affected since lack of knowledge, inadequate regulation, and outdoor air pollution lead to severe daily exposure to pollutants. However, only a limited number of studies have attempted to understand how indoor air pollution affects developing countries like India. To address this gap, we present spatiotemporal measurements of air quality from 30 indoor sites over six months during summer and winter seasons. The sites are geographically located across four regions of type: rural, suburban, and urban, covering the typical low to middle-income population in India. The dataset [^1] https://github.com/prasenjit52282/dalton-dataset (Access:August 17, 2024) contains various types of indoor environments (e.g., studio apartments, classrooms, research laboratories, food canteens, and residential households), and can provide the basis for data-driven learning model research aimed at coping with unique pollution patterns in developing countries. This unique dataset demands advanced data cleaning and imputation techniques for handling missing data due to power failure or network outages during data collection. Furthermore, through a simple speech-to-text application, we provide real-time indoor activity labels annotated by occupants. Therefore, environmentalists and ML enthusiasts can utilize this dataset to understand the complex patterns of the pollutants under different indoor activities, identify recurring sources of pollution, forecast exposure, improve floor plans and room structures of modern indoor designs, develop pollution-aware recommender systems, etc.
@@ 3.2 Annotating Activities of Daily Living
3.3 Measurement Sites & Demographics of the Occupants
@@
In total, there were 46 occupants in the measurement sites of the study. Among them, 24 occupants actively participated by providing the annotations for activity and event context of their indoor space using the vocalAnnot application as discussed in Sec.  3.2. For instance, in a household with four occupants, two of them annotated indoor activities and events for everyone. The occupants include college students, university staff, professors, homemakers, canteen owners, etc. Among them, 27 were male, and 19 were female. Table 2 depicts the female participation ratio and the socioeconomic background of the occupants. The majority of the occupants are between 20 and 40 years old. Thus, they are familiar with using Android apps during their everyday routines. Nevertheless, the youngest is a three-year-old child, while the oldest is an 84-year-old adult. Individuals, either too young or too old, do not actively participate in this study, and other family members report their indoor activities.
@@ 3.4 License and Consent
Figure 4: Split air conditioners suffer from compromised ventilation compared to legacy window air conditioners simply to improve power efficiency. The figure clearly depicts the accumulation of VOC and CO2 over time when split AC is on. Meanwhile, we observe consistent ventilation when the window AC is on, achieving healthy air quality with time. The relative humidity is slightly higher for window AC (i.e., within the comfort range, 30–50%).
@@ 5 Limitations and Conclusion
To conclude, in this work, we have collected a cross-sectional indoor air quality dataset from low to middle-income households and indoor sites in India. Unlike publicly available data, our dataset captures the indoor pollution dynamics of developing countries. Data was collected over six months across 30 indoor sites, including research labs, canteens, classrooms, studio apartments, and residential properties, with 46 occupants, among which 24 actively participated by providing annotations on indoor activities. Additionally, we include the geometry of the measurement sites, which can help understand the impact of floor plans on pollutant spread. We believe that this dataset can aid environmentalists, architects, and ML researchers in improving indoor designs, addressing pollution challenges, and developing models for predicting indoor pollution patterns.
@@ A.2 Composition
The dataset presents spatiotemporal air quality measurements from 30 indoor sites (e.g., studio apartments, classrooms, research laboratories, food canteens, and residential households) over six months during the summer and winter seasons (89.1M samples, totaling 13646 hours of air quality data and 3957 activity annotations from 24 participants among total 46 occupants).
@@ A.5 Preprocessing/cleaning/labeling
The dataset is cleaned and organized with the processing pipeline shown in Figure 8. First, the replicas from the log servers are merged, and any duplicates are dropped. Second, the sensor readings are validated based on the reported range of each sensor while marking any missing data segments. Next, missing values are imputed with the nearest fill method if the segment is less than 15 mins. Further, the data is smoothed with a 60-second rolling window to reduce noise before imputing remaining missing values with zero. Finally, the change-points in the sensor readings are marked to denote potential pollution events in the indoor space. After preprocessing, three new columns are computed from the sensor readings, as shown in the figure. The utility of the derived columns is as follows:
@@
Figure 9: Annotation processing pipeline.
We processed the raw annotations in five steps as shown in Figure 9. First, duplicates are removed, extra spaces are trimmed, and text is converted to lowercase. Second, we replace person names with a _NAME_ token to protect identities. Next, combined annotations are segregated into semantically valid segments as shown in the example in Figure 9. As a fourth step, we perform time synchronization for the split annotations based on their description. Finally, we correct any misspelled words resulting from incorrect speech-to-text conversion (e.g., frying can be annotated as flying or crying) in the annotation Android app, ensuring annotation accuracy and usability of the dataset.
@@ A.8 Maintenance
Figure 10: Daily indoor trends by week and month. We observe higher concentrations of VOC and CO2 during the winter months. The figure is reproduced here and depicted in work [54].
@@ Appendix B More Dataset Information
Figure 11: Missing data in the dataset.
B.1 Missing Data
Figure 11 illustrates the distribution of missing data percentage in each file within the dataset. We observe that missing data has an 11.82% mean with a standard deviation of 26.4%. However, the median and mode of missing data distribution are 0.0%, indicating high skewness of the distribution and assuring that most of the files have minimal missing data samples, as shown with a higher density of scatter points on the left side of the figure. The maximum missing data in the files is 97.91%. The boxplot reveals that most files have missing data percentages within the 0-10% range, implying the dataset’s quality.
B.2 Application of the Dataset
B.2.1 Analysis: Typical Pollution Sources & Spread of Pollutants
The dataset provides unique opportunities to analyze the spread of pollutants from typical sources like cooking, indoor gathering, eating, etc., due to multi-device deployment and the contextual activity annotations provided by the participants in the sites. Our prior study [54] explores such aspects of the dataset and shows the spreading, long-term lingering, and trapping of pollutants in various floor plans, room structures, and airflow patterns. Please refer to the study for more details.
(a) Activities in research lab R1
(b) Frequently cooked Food items in H1
Figure 12: Applications of the dataset and collected data from two sites R1, H1 - (a) detection of indoor activities, (b) Identification of cooked food items from pollution patterns.
(a) RF: Activity
(b) RF: Food
(c) NN: Activity
(d) NN: Food
Figure 13: Confusion Matrix of indoor activity and food item classification with (a)(b) Random Forest with max estimators 50 and max depth 10, (c)(d) Neural Network with three 64-neuron hidden layer.
B.2.2 ML Application: Identification of Indoor Activities
The dataset can be utilized to understand fluctuations in indoor pollutants (i.e., carbon dioxide, volatile organic compounds, particulate matter) and environmental parameters (i.e., temperature, humidity) that are influenced by indoor activities. We have considered eight indoor activities in a research lab (R1) that can be grouped into three categories based on monitoring motive: (i) Engagement and occupancy (i.e., enter, exit), (ii) Occupant behavior (i.e., fan on/off, AC on/off), (iii) Prohibited practices (i.e., gathering, eating). The class distribution is shown in Figure 12a.
Table 4 summarizes the detailed evaluation of 70-30 random split and 5-fold cross-validation experiments across seven machine learning models with varying parameters. As shown in Figure 13a and Figure 13c, we observe that random forest and neural network misclassify when someone enters or exits the lab. Moreover, exiting and turning on the fan is also confusing for the models. The degree of such confusion increases as we go for simpler models as per Table 4. As we evaluate for even simpler models like logistic regression and naive bayes, the overall performance significantly degrades to 76.4% and 39.9% F1-score, respectively. The primary reason is interleaved activities performed by the lab members, where the pollution signatures get convoluted due to multiple factors. For instance, a member can enter the lab and turn on the fan, resulting in a mixed influence on the pollution data. Moreover, the lab protocol also plays a role by insisting that the members turn on the fans when they leave the lab. Our benchmarking found that the best-performing model, random forest, shows 97.7% testing F1-score. Please refer to our work [55] for more details.
B.2.3 ML Application: Detection of Cooked Food Items for Automated Food Journalling
The dataset also shows potential applications for food item classification from the pollution signatures captured in the kitchen of household H1. We have considered eleven frequently cooked foods (i.e., ladies finger, lentils, egg, fish, poppy seeds, potato, pointed gourd, rice, roti, leafy vegetables, tea). The class distribution is shown in Figure 12b. With similar features computed in our work [55], we reported our benchmarking on seven machine learning algorithms in Table 5 for classifying the food items. As per the table, random forest and neural network perform best with 98.3% and 99.5% testing F1-score in 70-30 split experiments (see the confusion matrix in Figure 13b and Figure 13d). However, in 5-fold cross-validation, the performance drops due to class imbalance and fewer cooking instances for only using H1 (as cooking generally occurs only twice a day).
As discussed in this paper, our benchmarking indicates the dataset’s utility for several downstream ML applications using pollution data from multiple sensors and available activity annotations.
Table 4: Performance of the ML models in 70-30 random split and 5-fold cross-validation experiments in activity detection.
Model Parameters 70-30 Random Split 5-Fold Cross-validation
Training (Weighted) Testing (Weighted) Accuracy (Mean) Accuracy (Std)
F1-score Precision Recall F1-score Precision Recall Train Test Train Test
SVM Linear kernel 0.831 0.837 0.833 0.817 0.827 0.821 0.495 0.497 0.0156 0.0176
Polynomial kernel 0.892 0.894 0.892 0.879 0.881 0.879 0.543 0.542 0.0041 0.0075
RBF kernel 0.815 0.836 0.82 0.798 0.82 0.804 0.53 0.53 0.0036 0.01
Naive Bayes Gaussian 0.403 0.727 0.399 0.399 0.717 0.39 0.424 0.419 0.0105 0.0047
Decision Tree Max depth10 0.949 0.95 0.949 0.924 0.924 0.924 0.967 0.951 0.0025 0.0051
Max depth 20 0.992 0.992 0.992 0.975 0.975 0.976 0.992 0.974 0.0008 0.0039
Max depth 30 0.992 0.992 0.992 0.976 0.976 0.976 0.992 0.975 0.0007 0.0039
Max depth 40 0.992 0.992 0.992 0.976 0.976 0.976 0.992 0.975 0.0007 0.0039
k-Nearest Neighbour Neighbour 10 0.981 0.981 0.981 0.975 0.975 0.975 0.985 0.979 0.0008 0.0022
Neighbour 20 0.972 0.972 0.972 0.967 0.967 0.967 0.983 0.979 0.0002 0.0024
Neighbour 30 0.959 0.96 0.96 0.956 0.956 0.957 0.979 0.976 0.0015 0.0044
Neighbour 40 0.946 0.946 0.947 0.947 0.947 0.947 0.975 0.972 0.0021 0.0036
Logistic Regression 0.791 0.801 0.796 0.764 0.777 0.77 0.581 0.577 0.0088 0.0145
Random Forest
Max estimator 30
Max depth 10
0.988 0.988 0.988 0.979 0.979 0.979 0.989 0.977 0.0005 0.0021
Max estimator 50
Max depth 10
0.988 0.988 0.988 0.977 0.977 0.977 0.989 0.979 0.0006 0.0049
Max estimator 100
Max depth 10
0.99 0.99 0.99 0.979 0.979 0.979 0.989 0.977 0.0012 0.0037
Neural Network Hidden [64, 64] 0.981 0.981 0.981 0.973 0.973 0.973 0.925 0.92 0.0229 0.0165
Hidden [64, 64, 64] 0.982 0.983 0.982 0.978 0.979 0.978 0.947 0.943 0.0104 0.0135
Hidden [128, 128] 0.981 0.982 0.981 0.979 0.979 0.979 0.912 0.91 0.0116 0.0114
Hidden [128, 128, 128] 0.982 0.982 0.982 0.978 0.978 0.978 0.95 0.946 0.0174 0.0203
Table 5: Performance of the ML models in 70-30 random split and 5-fold cross-validation experiments in food item detection.
model Parameters 70-30 Random Split 5-Fold Cross-validation
Training (Weighted) Testing (Weighted) Accuracy (Mean) Accuracy (Std)
F1-score Precision Recall F1-score Precision Recall Train Test Train Test
SVM Linear kernel 0.741 0.781 0.754 0.734 0.781 0.747 0.681 0.33 0.064 0.1041
Polynomial kernel 0.865 0.895 0.868 0.866 0.897 0.87 0.601 0.426 0.0133 0.0592
RBF kernel 0.838 0.873 0.842 0.837 0.871 0.841 0.625 0.419 0.0134 0.0809
Naive Bayes Gaussian 0.334 0.526 0.318 0.334 0.516 0.317 0.363 0.223 0.0206 0.0242
Decision Tree Max depth10 0.974 0.975 0.974 0.961 0.962 0.961 0.941 0.387 0.0328 0.064
Max depth20 0.999 0.999 0.999 0.984 0.984 0.984 0.999 0.374 0.0003 0.0634
Max depth30 0.999 0.999 0.999 0.984 0.984 0.984 0.999 0.374 0.0003 0.0634
Max depth40 0.999 0.999 0.999 0.984 0.984 0.984 0.999 0.374 0.0003 0.0634
k-Nearest Neighbour Neighbour 10 0.991 0.991 0.991 0.982 0.982 0.982 0.965 0.407 0.0047 0.0885
Neighbour 20 0.963 0.964 0.963 0.948 0.949 0.948 0.921 0.388 0.008 0.0836
Neighbour 30 0.921 0.922 0.921 0.912 0.914 0.912 0.864 0.386 0.0112 0.0732
Neighbour 40 0.847 0.851 0.849 0.849 0.857 0.852 0.823 0.368 0.0172 0.0724
Logistic Regression 0.659 0.695 0.686 0.652 0.692 0.679 0.603 0.415 0.0231 0.0674
Random Forest
Max estimator 30
Max depth 10
0.988 0.988 0.988 0.978 0.98 0.979 0.995 0.521 0.0033 0.1361
Max estimator 50
Max depth 10
0.991 0.991 0.991 0.982 0.983 0.982 0.993 0.532 0.0056 0.1298
Max estimator 100
Max depth 10
0.997 0.997 0.997 0.994 0.994 0.994 0.996 0.524 0.0031 0.1234
Neural Network Hidden [64, 64] 0.99 0.99 0.989 0.983 0.983 0.983 0.799 0.368 0.057 0.0539
Hidden [64, 64, 64] 0.998 0.998 0.998 0.996 0.996 0.996 0.913 0.42 0.0522 0.0928
Hidden [128, 128] 0.999 0.999 0.999 0.994 0.994 0.994 0.892 0.434 0.0385 0.1216
Hidden [128, 128, 128] 0.997 0.997 0.997 0.995 0.995 0.995 0.965 0.443 0.0226 0.1446
B.3 Additional Insights
The Figure 14a and Figure 14b present the spatiotemporal spread of VOC and CO2 in H1 during and after cooking activity in Kitchen (A). Cooking starts at $t$ time and gets over by $t$+11 mins. We observe that VOC is hard to ventilate compared to CO2. VOC spreads even after 6 minutes (i.e., $t$+17 mins), whereas CO2 depletes to normal levels. In Figure 14a (subfigure $t$+48 mins), we observe trapping of VOC in the bedroom (E).
The Figure 15a and Figure 15b shows the spread of VOC and CO2 in H4 during and after cooking activity in Kitchen (A). Cooking ended at $t$+33 mins. Unlike H1, we observe increased levels of VOC and CO2 even after 14 mins (i.e., Figure 15, subfigures $t$+48 mins). Moreover, long-term lingering and trapping of the VOC can be observed in Figure 15a (subfigure $t$+90 mins) due to the congested room structure of H4.
Figure 14: Spatiotemporal spread of – (a) VOC, and (b) CO2 from the kitchen in Household H1. The figure is reproduced to provide insights on the dataset and depicted in work [54].
Figure 15: Spatiotemporal spread of – (a) VOC, and (b) CO2 from the kitchen in Household H4.The figure is reproduced to provide insights on the dataset and depicted in work [54].
@@ B.3 Additional Insights
Site ID (Lat,Long) Devices #Devices Site Area (Sqft) Floor Plan #F/#M Duration (Hrs) #Samples Activity Annot. (Hrs) Occupants
@@ 95 Kitchen
A1
(22.316371948187594,
87.30576795184774)
101 Study Desk 1 150 1/0 146 226888 48 P28
A2
((22.30971013354728,
87.30784886676756))
105 Study Desk 1 150 0/1 289 193557 P29
A3
(22.30937218266447,
87.31064479668473)
109 Study Desk 1 180 0/1 344 1098827 96 P30
A4
(22.318929727398313,
87.30650918833601)
120 Study Desk 1 150 1/0 125 384975 P31
A5
22.318929727398313,
87.30650918833601
121 Study Desk 1 150 1/0 1 77 1 P32
A6
(22.31648291562444,
87.29538496948412)
122 Study Desk 1 100 0/1 51 154398 51 P33
A7
(22.30971013354728,
87.30784886676756)
113 On bed 1 150 0/1 55 54741 55 P34
A8
(22.30971013354728,
87.30784886676756)
151 Study Desk 1 150 0/1 60 189141 P35
@@ 52 Desk2
R2
(22.3179564762161,
87.30920753796912)
102 Desk Lab 319 1 320 2/2 367 1161570 72 P43
R3
(22.3179564762161,
87.30920753796912)
108 Left Desk 1 616 0/1 243 750745 72 P44
@@ 71 Desk1
F1
(22.30936287704651,
87.31055134200086)
23 Kitchen 1 150 2/0 450 631193 P46
F2
(22.32048015405153,
87.30800205041001)
24 Kitchen 1 150 450 631193
C1
(22.3179564762161,
87.30920753796912)
25 Teacher Desk 1 500 333 590272
C2
(22.3179564762161,
87.30920753796912)
26 Teacher Desk 1 500 53 158256